solve the fibinacci recursion formula to a closed form usung the inverse z-transform
Z --U Fix is furtf 6 F-1 = f -z = 1 fica fkul + F 1 U = Z CU U 42
Answer #1

form using ! for mula to a closed solve Fibonacci recursion the inverse a-tromsform. . . So Fibonacci recurrence ...Fk+2 = FkZC Uk+2) = z2(UZ) - uro) - CU) ) = za ufz)- 02?12 = Z²uEl-Z ZC Uk +1) = Z CUZ) - 410)) = ZUZ) - 0. z cuk+i)= ZUZ) z cuk) -ua) Z? 7-1 - +1 +1 +4 22 2. (1755). op 12-1995) 12–195) -12) (13 Da 03 zcantina za Apply Inversezat zlewmeyin) - Z Zurz IC Z